S.S. d.f. M.S. F
Between 6.892
Within 3146 121
TOTAL 28

Use the =FDIST() function in Excel to locate the p-value for this ANOVA:
p-value =
Report answer accurate to at least 4 decimal places.
Answer #1

From given ANOVA we get,

within d.f. = SS within / MS within = 3146/121 = 26

Between d.f. = Total d.f. - within d.f. = 28 - 26 = 2

Now, we find, p-value using FDIST()

= FDIST(6.892, 2, 26)

= 0.0040

=> p-value = 0.0040